# Association of the IceCube neutrinos with CAZ blazar light curves

### v1

@@ 2.2.3 CAZ light curves and variability
In K25-Cat we compiled as long and as high cadence optical light curves as possible for the 7918 sources in the CAZ catalog. We extracted multiband light curves from the CRTS777https://crts.caltech.edu/ (e.g., Drake et al. 2009), ATLAS888https://fallingstar-data.com/ (e.g., Tonry et al. 2018), and ZTF999https://ztf.caltech.edu/ (e.g., Bellm et al. 2019) all-sky surveys. This is done by performing forced-photometry on the coordinates of the CAZ sources. We additionally took data from the Tuorla blazar monitoring program101010https://tuorlablazar.utu.fi/ (e.g., Nilsson et al. 2018) and the Katzman Automatic Imaging Telescope (KAIT111111https://w.astro.berkeley.edu/bait/kait.html) monitoring program (e.g., Filippenko et al. 2001; Cohen et al. 2014), when possible.
@@
To ensure maximal cadence for the CAZ light curves, we merged the filters of the different surveys121212There are 11 filter+survey combinations in total: one from CRTS (no filer, but closest is V-band), five from ATLAS (o, c, r, g, and i), three from ZTF (r, g, and i), one from Tuorla (R-band), and one from KAIT (V-band). by translating them in flux density space. This was done, in order of decreasing priority, by minimizing the difference between simultaneous flux density measurements of two filters (within 1, or then days), or minimizing the difference between the average flux density of two filters (within their overlap periods, or then within a 50-day extension). The procedure to obtain the final CAZ light curves is described in more detail in Sect. 3 of K25-Cat. We note that all CAZ light curves are publicly available. In total, there are 7722 CAZ light curves with 10 data points and 30 d duration.
@@
In K25-Cat we characterized the variability of the CAZ light curves, and then determined periods of enhanced emission and prominent flaring periods in the variable ones. These periods are used in this study to search for temporal coincidences with neutrinos. In short, we focused on the CRTS and ZTF data, while ignoring ATLAS data due to its noisiness, to determine whether a light curve is confidently variable. We quantified their variability using fractional variability (; e.g., Sokolovsky et al. 2017) and used white dwarf light curves as control. In short, a CAZ light curve is considered variable if either its CRTS or its ZTF reach certain critical thresholds. For more details on the variability quantification, we refer the reader to Sect. 4.1 of K25-Cat. Of the 7722 dense and long enough CAZ light curves, 2798 (36.2%) were determined to be confidently variable. Notably, this does not mean that the other 4924 (63.8%) light curves are confidently non-variable, but that their variability is comparable to that of white dwarfs within the predefined thresholds. Since a typical CAZ light curve can be rather noisy, the threshold to confidently “detect” variability is rather high.
@@
Since in the main analysis of this study (Sect. 3) we only perform spatio-temporal correlation tests, we solely focus on the variable blazars whose flares can be confidently identified. This inadvertently excludes 63.7% of RFC† and 49.7% of 4LAC blazars from our analysis131313This exclusion rate was set at a similar fraction (i.e., 70%) in the simulations of K25-Sim.. We provide the parameter availability of the variable RFC† and 4LAC samples in Table 1 columns (3) and (5), respectively. Nonetheless, we note that, in Sect. 4.7, we perform a complementary, spatial-only analysis which uses all RFC† and 4LAC blazars regardless of their variability.
@@ 2.2.4 Periods of enhanced emission (BB95 periods)
In K25-Cat, to characterize periods of enhanced emission and prominent flaring (Sect. 2.2.5), we applied the Bayesian block (BB; Scargle et al. 2013) algorithm to the CAZ light curves. Using inherently non-variable white dwarf light curves, which were constructed identically to blazar light curves, we calibrated the BB algorithm to partially account for noise-driven variability. Furthermore, when constructing the BB light curves, we marked extended periods (60 d) without data as seasonal gaps141414Since optical observations are only possible during dark skies, long-term optical light curves often exhibit extended periods of gaps in the data due to the seasonal inability to observe the sources..
Subsequently, we defined periods of enhanced emission as those BB whose averaged brightness exceeds the 95th percentile flux density level. These are referred to as “BB95”. Such strict flux density condition ensures that BB95 periods exclusively trace the most extreme optical periods of a variable light curve. As shown in Fig. 7 of K25-Cat, BB95 durations have an extended range, with typical values of 2–13 d. Unsurprisingly, almost all variable CAZ light curves result in some BB95: 1161 (99.2%) of 1170 variable RFC† sources, and 1906 (99.4%) of 1917 variable 4LAC sources. The few light curves which do not result in any BB95 have such large flux density error bars that none of their binned flux densities reach the 95th percentile level.
While BB95 are guaranteed to characterize the most extreme activity periods of a blazar, they suffer from two main caveats: (1) they are easily affected by extreme outliers; and (2) they miss most flaring periods since only a handful of flares reach the most extreme flux densities. As an alternative to BB95, in the following section (2.2.5) we describe how prominent flaring periods are identified in variable CAZ light curves.
@@ 2.2.5 Prominent flaring periods (BBHOP flares)
In K25-Cat, to determine prominent flaring periods, we initially applied the BB algorithm to the CAZ light curves. Then we searched for BB whose average flux density corresponds to a local maximum or minimum (i.e., “hills” and “valleys” in BB flux density, respectively). Subsequently, using a hopping algorithm (e.g., HOP, Eisenstein & Hut 1998), we determined valley-hill-valley groups without going through any seasonal gap. These BB groups are referred to as “BBHOP” flares (e.g., Meyer et al. 2019). To ensure that major flaring periods are characterized in their entirety and that small flux density variations do not result in premature BBHOP flares, we iteratively merged all pairs of neighboring BBHOP flares if: (1) their rise or fall time was short (2 d); or (2) their rise or fall flux density significance change was small (3). For more details, see Sect. 4.3 of K25-Cat.
Merged BBHOP flares (hereafter simply referred to as BBHOP flares) characterize most flaring periods in reasonably high-cadence portions of a light curve, even when the flux rise is relatively small. However, we were interested in the most prominent flaring periods, that is, those with the most prominent rise in flux. Therefore, we defined prominent BBHOP flares as those that: (1) occurred in a variable CAZ light curve; (2) had a peak BB brighter than the 75th percentile flux density; (3) had 3 rise or fall flux density significance change; and (4) did not consist of three BBs only with just one data point in their peak BB. These conditions resulted in 25194 prominent BBHOP flares across 2639 variable CAZ light curves. In terms of the blazar samples used in this study, 1108 (94.7%) of 1170 variable RFC† sources and 1872 (97.7%) of 1917 variable 4LAC sources have at least one prominent BBHOP flare. Lastly, we note that prominent BBHOP flares have a wide range of total durations, with typical values of 30–130 d.
While prominent BBHOP flares characterize most major flaring periods, they suffer from a couple of caveats. Firstly, the efficiency of the BBHOP algorithm to characterize flares depends strongly on the cadence and continuity of the data. When the cadence drops or when there are seasonal gaps in the data, it is not possible to reliably identify local minima and maxima in flux density. Secondly, outliers can result in BBHOP flares. Unfortunately, low cadence, seasonal gaps, and noisy data are prevalent in the CAZ light curves, despite our attempts to mitigate them (for more details, see K25-Cat).
@@ 2.2.6 BB95 periods as the peak of BBHOP flares
Since both BB95 and prominent BBHOP flares have strengths and weaknesses for characterizing outburst periods in blazars, we opt to use both of them. Additionally, we consider them simultaneously to identify the peak of the most extreme flares. So, as a third temporal metric in our spatio-temporal correlation tests, we look for BB95 periods which form the peak of a prominent BBHOP flare. Notably, this combined metric does not probe for BB95 periods which occur in the rise or decay segments of a BBHOP flare, but only those which coincide with its peak. For example, in the association of the blazar CAZJ0207+0950 with the neutrino IC190629A (Fig. 4), both the BB95 and BBHOP metrics are met separately, but the combined metric is not met. By identifying the peaks of the most prominent flaring periods via two criteria, we reduce the chance of misidentifying noise-driven scatter as intrinsic. The obvious downside of this combined metric is its need for a high cadence.
@@ 3 Analysis
In this paper we search for spatio-temporal correlation between optical light curves of blazars and IceCube neutrinos. The spatio-temporal test is similar to the one we used in H21 and K24.
@@
As a first step, we identify which blazars are spatially associated with the neutrino events. We do this by searching for blazar coordinates that fall within the enlarged (by 1∘ added in quadrature) 90%-likelihood error region of the neutrinos. The enlargement is introduced to account for potential unknown systematic errors in the spatial reconstruction of the events (e.g., Aartsen et al. 2013; see Sect. 3.3 of K24 for more details). Additionally, previous works (e.g., Plavin et al. 2020, H21, K24, Abbasi et al. 2023a) have shown that such an enlargement is necessary to obtain any hint of a correlation with blazars, since several spatio-temporally associated blazars narrowly miss the published 90%-likelihood error region of some well-reconstructed neutrino events. We emphasize that our spatio-temporal method, specifically with a top-hat-like weighting scheme, is robust to such error region enlargements. In fact, our simulations in K25-Sim (see Sect. 4.2.6 therein) showed that enlarging the neutrino error regions yields conservative estimates for the strength of the underlying blazar-neutrino correlation. Therefore, in this paper, we only perform our spatio-temporal tests using the enlarged 90%-likelihood error regions.
Once the spatially associated blazars are identified, we search their CAZ light curves and determine whether they are temporally associated with their spatially associated neutrino. A temporal association is established if the neutrino arrived within 2 d of a BB95 (Sect. 2.2.4), during a prominent BBHOP flare (Sect. 2.2.5), or within 2 d of a BB95 which is the peak of a prominent BBHOP flare (Sect. 2.2.6). We allow for a 2 d time window between a BB95 and the neutrino arrival time, because 2 d was the maximum limit of quasi-simultaneity used throughout K25-Cat. Namely, we used it when merging the filters onto each other (Sect. 2.2.3) and when merging neighboring BBHOP flares together (Sect. 2.2.5). We emphasize that, by construction, temporal associations only occur with variable blazars.
@@
We then count the spatio-temporal associations to construct our test-statistic (TS) parameter. To obtain the global strength of the blazar-neutrino spatio-temporal correlation, we determine the probability of obtaining the observed TS parameter by chance via a Monte Carlo simulation. We randomize the blazar RA coordinates151515Randomization in declination is avoided since the reliability of the reconstruction of the neutrino events is strongly declination dependent. and then repeat the above process to count the number of random spatio-temporal associations, resulting in a random TS parameter. This randomization is repeated times such that we obtain random TS parameters. Subsequently, using the following equation we obtain the global strength of the blazar-neutrino spatio-temporal correlation:
@@
where is the number of random TS parameters greater than or equal to the observed TS parameter, and is the total number of random TS parameters (Davison & Hinkley 1997); in this work . This process is visualized in Fig. 3 of K24.
@@
We note that, in H21 and K24, we constructed the TS parameters via both averaging and counting techniques. However, in this work, we only focus on the counting technique since our extensive simulations in K25-Sim showed that counting offers a greater detection power than averaging. Additionally, K25-Sim showed that employing a weighting scheme to account for poorly reconstructed neutrino events is better than setting neutrino sample cuts (e.g., or cuts). Among the tested weighting schemes, the top-hat scheme performed better than the Gaussian one. Therefore, before constructing the TS parameter, we weigh them down using a top-hat weight (; first introduced in K24). So, the global TS parameter is equal to where is the top-hat weight of the spatio-temporal association and is the total number of spatio-temporal associations globally.
@@
where is the global median of (see Eqn. 1), the distance between the center and the edge of the error region at a given phase angle , and the distance between the blazar and the center of the error region along (see Fig. 1 of K25-Sim). As noted in Sect. 2.1, for the published error regions. However, here we use the enlarged error regions, whose larger result in . When applied to all 356 enlarged neutrino events of the updated IceCat1+, the maximum, median, and minimum of are 0.997, 0.316, and 0.003, respectively.
Notably, reduces the effect of poorly localized events (i.e., those with ) on the statistics, while preventing the well-localized events (i.e., those with ) from dominating the statistics. This makes the top-hat weighting scheme superior to Gaussian weighting schemes, which are typically dominated by associations made with the most highly localized events. This is a caveat for the Gaussian weights since mid-range events (i.e., those with and ) are more numerous than the most highly localized events, even after accounting for their smaller (see Sect. 4.2.5 of K25-Sim). The advantage of comes from its ability to recover more signal from the mid-range events.
We note that, due to the statistical nature of our analysis, not all of the identified associations are expected to be real. Following our previous works (H21, K24, and K25-Sim), the global TS parameter is constructed by counting the (weighted) number of unique spatio-temporally associated blazar-neutrino pairs. This allows for: (1) one blazar to be associated with multiple neutrinos; and (2) one neutrino to be associated with multiple flaring blazars. Notably, the latter is physically impossible. Yet, it enhances the statistical power of the test, because a neutrino associated with multiple flaring blazars is more likely to have come from a blazar than a neutrino associated with only one flaring blazar. Therefore, our global TS parameters should be taken as statistical measures rather than real counts.

### v2

@@ 2.2.3 CAZ light curves and variability
In K26 we compiled as long and as high cadence optical light curves as possible for the 7918 sources in the CAZ catalog. We extracted multiband light curves from the CRTS777https://crts.caltech.edu/ (e.g., Drake et al. 2009), ATLAS888https://fallingstar-data.com/ (e.g., Tonry et al. 2018), and ZTF999https://ztf.caltech.edu/ (e.g., Bellm et al. 2019) all-sky surveys. This was done by performing forced-photometry on the coordinates of the CAZ sources. We additionally took data from the Tuorla blazar monitoring program101010https://tuorlablazar.utu.fi/ (e.g., Nilsson et al. 2018) and the Katzman Automatic Imaging Telescope (KAIT111111https://w.astro.berkeley.edu/bait/kait.html) monitoring program (e.g., Filippenko et al. 2001; Cohen et al. 2014), when possible.
@@
To ensure maximal cadence for the CAZ light curves, we merged the filters of the different surveys121212There are 11 filter+survey combinations in total: one from CRTS (no filer, but closest is V-band), five from ATLAS (o, c, r, g, and i), three from ZTF (r, g, and i), one from Tuorla (R-band), and one from KAIT (V-band). by translating them in flux density space. This was done, in order of decreasing priority, by minimizing the difference between simultaneous flux density measurements of two filters (within 1, or then days), or minimizing the difference between the average flux density of two filters (within their overlap periods, or then within a 50-day extension). The procedure to obtain the final CAZ light curves is described in more detail in Sect. 3 of K26. We note that all CAZ light curves are publicly available. In total, there are 7722 CAZ light curves with 10 data points and 30 d duration.
@@
In K26, we characterized the variability of the CAZ light curves, and then determined periods of enhanced emission and prominent flaring periods in the variable ones. These periods are used in this study to search for temporal coincidences with neutrinos. In short, we focused on the CRTS and ZTF data, while ignoring ATLAS data due to its noisiness, to determine whether a light curve is confidently variable. We quantified their variability using fractional variability (; e.g., Sokolovsky et al. 2017) and used white dwarf light curves as control. A CAZ light curve is considered variable if either its CRTS or its ZTF reach certain critical thresholds. For more details on the variability quantification, we refer the reader to Sect. 4.1 of K26. Of the 7722 dense and long enough CAZ light curves, 2798 (36.2%) were determined to be confidently variable. Notably, this does not mean that the other 4924 (63.8%) light curves are confidently non-variable, but that their variability is comparable to that of white dwarfs within the predefined thresholds. Since a typical CAZ light curve can be rather noisy, the threshold to confidently “detect” variability is rather high.
@@
Since for the main analysis of this study we only performed spatiotemporal correlation tests (requiring flares; see Sect. 3), here we solely focus on variable blazars whose flares can be confidently identified. This inadvertently excludes 63.7% of RFC† and 49.7% of 4LAC blazars from our analysis131313This exclusion rate was set at a similar fraction (i.e., 70%) in the simulations of K25.. We provide the parameter availability of the variable RFC† and 4LAC samples in Table 1, columns (3) and (5), respectively. Nonetheless, we note that, in Sect. 4.7, we describe our complementary, spatial-only analysis, which uses all RFC† and 4LAC blazars regardless of their variability.
@@ 2.2.4 Periods of enhanced emission (BB95 periods)
In K26, to characterize periods of enhanced emission and prominent flaring (Sect. 2.2.5), we applied the Bayesian block (BB; Scargle et al. 2013) algorithm to the CAZ light curves. Using inherently non-variable white dwarf light curves, which were constructed identically to blazar light curves, we calibrated the BB algorithm to partially account for noise-driven variability. Furthermore, when constructing the BB light curves, we marked extended periods (60 d) without data as seasonal gaps141414Since optical observations are only possible during dark skies, long-term optical light curves often exhibit extended periods of gaps in the data due to the seasonal inability to observe the sources..
Subsequently, we defined periods of enhanced emission as those BB whose averaged brightness exceeds the 95th percentile flux density level, referred to as “BB95.” Such strict flux density condition ensures that BB95 periods exclusively trace the most extreme optical periods of a variable light curve. As shown in Fig. 7 of K26, BB95 durations have an extended range, with typical values of 2–13 d. Unsurprisingly, almost all variable CAZ light curves result in some BB95: 1161 (99.2%) of 1170 variable RFC† sources, and 1906 (99.4%) of 1917 variable 4LAC sources. The few light curves that do not result in any BB95 have such large flux density error bars that none of their binned flux densities reach the 95th percentile level.
While BB95 are guaranteed to characterize the most extreme activity periods of a blazar, they suffer from two main caveats: (1) they are easily affected by extreme outliers and (2) they miss most flaring periods since only a handful of flares reach the most extreme flux densities. As an alternative to BB95, in the following subsection, we describe how prominent flaring periods are identified in variable CAZ light curves.
@@ 2.2.5 Prominent flaring periods (BBHOP flares)
In K26, to determine prominent flaring periods, we initially applied the BB algorithm to the CAZ light curves. Then we searched for BB whose average flux density corresponds to a local maximum or minimum (i.e., “hills” and “valleys” in BB flux density, respectively). Subsequently, using a hopping algorithm (e.g., HOP, Eisenstein & Hut 1998), we determined valley-hill-valley groups without going through any seasonal gap. These BB groups are referred to as “BBHOP” flares (e.g., Meyer et al. 2019). To ensure that major flaring periods are characterized in their entirety and that small flux density variations do not result in premature BBHOP flares, we iteratively merged all pairs of neighboring BBHOP flares if: (1) their rise or fall time was short (2 d); or (2) their rise or fall flux density significance change was small (3). For more details, see Sect. 4.3 of K26.
Merged BBHOP flares (hereafter, simply referred to as BBHOP flares) characterize most flaring periods in reasonably high-cadence portions of a light curve, even when the flux rise is relatively small. However, we were interested in the most prominent flaring periods, that is, those with the most prominent rise in flux. Therefore, we defined prominent BBHOP flares as those that: (1) occurred in a variable CAZ light curve; (2) had a peak BB brighter than the 75th percentile flux density; (3) had 3 rise or fall flux density significance change; and (4) did not consist of three BBs only with just one data point in their peak BB. These conditions resulted in 25194 prominent BBHOP flares across 2639 variable CAZ light curves. In terms of the blazar samples used in this study, 1108 (94.7%) of 1170 variable RFC† sources and 1872 (97.7%) of 1917 variable 4LAC sources have at least one prominent BBHOP flare. Lastly, we note that prominent BBHOP flares have a wide range of total durations, with typical values of 30–130 d.
While prominent BBHOP flares characterize most major flaring periods, they suffer from a couple of caveats. Firstly, the efficiency of the BBHOP algorithm to characterize flares depends strongly on the cadence and continuity of the data. When the cadence drops or when there are seasonal gaps in the data, it is not possible to reliably identify local minima and maxima in flux density. Secondly, outliers can result in BBHOP flares. Unfortunately, low cadence, seasonal gaps, and noisy data are prevalent in the CAZ light curves, despite our attempts to mitigate them (for more details, see K26).
@@ 2.2.6 BB95 periods as the peak of BBHOP flares
Since both BB95 and prominent BBHOP flares have strengths and weaknesses for characterizing outburst periods in blazars, we opt to use both of them. Additionally, we consider them simultaneously to identify the peak of the most extreme flares. So, as a third temporal metric in our spatiotemporal correlation tests, we look for BB95 periods which form the peak of a prominent BBHOP flare. Notably, this combined metric does not probe for BB95 periods which occur in the rise or decay segments of a BBHOP flare, but only those which coincide with its peak. For example, in the association of the blazar CAZJ0207+0950 with the neutrino IC190629A (Fig. 4), both the BB95 and BBHOP metrics are met separately, but the combined metric is not met. By identifying the peaks of the most prominent flaring periods via two criteria, we reduce the chance of misidentifying noise-driven scatter as intrinsic. The obvious downside of this combined metric is its need for a high cadence.
@@ 3 Analysis
In this work, we searched for spatiotemporal correlation between optical light curves of blazars and IceCube neutrinos. The spatiotemporal test we adopted is similar to the one we used in H21 and K24. Here, we use blazar flares identified in variable CAZ optical light curves in K26, which limits the blazar sample of the spatiotemporal search of this paper to a subset of optically variable blazars only.
@@
As a first step, we identified blazars that are spatially associated with the neutrino events. We did this by searching for blazar coordinates that fall within the enlarged (by 1∘ added in quadrature) 90%-likelihood error region of the neutrinos. The enlargement was introduced to account for potential unknown systematic errors in the spatial reconstruction of the events (e.g., Aartsen et al. 2013; see Sect. 3.3 of K24 for more details). Additionally, previous works (e.g., Plavin et al. 2020, H21, K24, Abbasi et al. 2023a) have shown that such an enlargement leads to a hint of a correlation with blazars, since several spatiotemporally associated blazars narrowly miss the published 90%-likelihood error region of some well-reconstructed neutrino events. We emphasize that our spatiotemporal method, specifically with a top-hat-like weighting scheme, is robust to such error region enlargements. In fact, our simulations in K25 (see Sect. 4.2.6 therein) showed that enlarging the neutrino error regions yields conservative estimates for the strength of the underlying blazar-neutrino correlation. Therefore, here we only performed our spatiotemporal tests using the enlarged 90%-likelihood error regions.
Once the spatially associated blazars are identified, we can search their CAZ light curves to determine whether they are temporally associated with their spatially associated neutrino. Based on the simple phenomenological assumption that an enhancement in neutrino production is accompanied by an enhancement in electromagnetic emission, we defined a temporal association based on the following Boolean condition: If the neutrino arrival time coincides with a CAZ optical flare, there is a temporal association; otherwise, there is no temporal association. This assumes simultaneity between neutrino and optical flares (within a few days, considering data limitations). In support of this scenario, there are theoretical models that predict optical flares lagging behind neutrino flares by 1 d (e.g., Rodrigues et al. 2025). However, we caution that there are also theoretical models which predict potential time delays of months to years between neutrino flares and their electromagnetic counterparts (e.g., Podlesnyi & Oikonomou 2025).
Using the CAZ light curves, we define a temporal association if the neutrino arrived within 2 d of a BB95 (Sect. 2.2.4), during a prominent BBHOP flare (Sect. 2.2.5), or within 2 d of a BB95, which is the peak of a prominent BBHOP flare (Sect. 2.2.6). Since the CAZ light curves suffer from numerous limitations (e.g., uneven sampling, seasonal gaps, outliers, potentially poorly color-corrected simultaneous data), quasi-simultaneity cannot be defined exactly and requires an error margin. Throughout K26, we used 2 d as the lower limit for such an error margin; namely, we used it when merging the filters onto each other (Sect. 2.2.3) and when merging neighboring BBHOP flares together (Sect. 2.2.5). This 2 d margin is especially necessary when checking for quasi-simultaneity with BB95 periods due to their rather short durations (e.g., as short as 1 d); however, it is not necessary for BBHOP flares because they typically constitute tens of BBs and are 30–130 d in duration.
@@
We then counted the spatiotemporal associations to construct our test-statistic (TS) parameter. To obtain the global strength of the blazar-neutrino spatiotemporal correlation, we determined the probability of obtaining the observed TS parameter by chance via a Monte Carlo simulation. We randomized the blazar RA coordinates151515Randomization in declination is avoided since the reliability of the reconstruction of the neutrino events is strongly declination dependent. and then repeated the above process to count the number of random spatiotemporal associations, resulting in a random TS parameter. This randomization was repeated times such that we obtain random TS parameters. Subsequently, using the following equation, we can obtain the global strength of the blazar-neutrino spatiotemporal correlation:
@@
where is the number of random TS parameters greater than or equal to the observed TS parameter, and is the total number of random TS parameters (Davison & Hinkley 1997); in this work, . This process is visualized in Fig. 3 of K24. Although this correlation strength is obtained through the uniform randomization of blazars in the sky, we verified that the results do not change if we exclude blazars within 10∘ from the Galactic plane.
@@
In H21 and K24, we constructed the TS parameters via both averaging and counting techniques. However, in this work, we only focus on the counting technique since our extensive simulations in K25 showed that counting offers a greater detection power than averaging. Additionally, K25 showed that employing a weighting scheme to account for poorly reconstructed neutrino events is better than setting neutrino sample cuts (e.g., or cuts). Among the tested weighting schemes, the top-hat scheme performed better than the Gaussian one. Therefore, before constructing the TS parameter, we weigh them down using a top-hat weight (; first introduced in K24). So, the global TS parameter is equal to where is the top-hat weight of the spatiotemporal association and is the total number of spatiotemporal associations globally.
@@
Here, is the global median of (see Eqn. 1), the distance between the center and the edge of the error region at a given phase angle, , and is the distance between the blazar and the center of the error region along (see Fig. 1 of K25). As noted in Sect. 2.1, for the published error regions. However, here, we use the enlarged error regions, whose larger values result in . When applied to all 356 enlarged neutrino events of the updated IceCat1+, the maximum, median, and minimum of are 0.997, 0.316, and 0.003, respectively.
Notably, reduces the effect of poorly localized events (i.e., those with ) on the statistics, while preventing the well-localized events (i.e., those with ) from dominating the statistics. This makes the top-hat weighting scheme superior to Gaussian weighting schemes, which are typically dominated by associations made with the most highly localized events. This is a caveat for the Gaussian weights since mid-range events (i.e., those with and ) are more numerous than the most highly localized events, even after accounting for their smaller (see Sect. 4.2.5 of K25). The advantage of comes from its ability to recover more signal from the mid-range events.
We note that due to the statistical nature of our analysis, not all of the identified associations are expected to be real. Following our previous works (H21, K24, and K25), the global TS parameter is constructed by counting the (weighted) number of unique spatiotemporally associated blazar-neutrino pairs. This allows for: (1) one blazar to be associated with multiple neutrinos; and (2) one neutrino to be associated with multiple flaring blazars. Notably, the latter is physically impossible. Yet, it enhances the statistical power of the test, because a neutrino associated with multiple flaring blazars is more likely to have come from a blazar than a neutrino associated with only one flaring blazar. Therefore, our global TS parameters should be taken as statistical measures rather than real counts. As a sanity check, we also test our spatiotemporal analysis by only considering one flaring blazar association for each neutrino event. In this approach, the TS parameter is calculated by counting the number of neutrino events which have at least one spatiotemporal association. This approach does not suffer from the second limitation mentioned above. We find that the correlation strengths remain largely identical to those obtained via the original formulation.